solve using elimination method
x+8y=31
2x+7y=35

Answer by mananth(16949)

1 x + 8 y = 31 .............1
2 x + 7 y = 35 .............2
Eliminate y
multiply (1)by -7
Multiply (2) by 8
-7x-56y =-217
16x+56y =280
Add the two equations
9x=63
/9
x=7
plug value of x in (1)
x+8y=31
7+8y=31
8y=31-7
8y=24
y=3